A population has mean and standard deviation of 20 and 3.15,

respectively. What is the standard error of the mean of the sampling

distribution in a sample size of 5? complete solution


1
Expert's answer
2021-05-07T09:59:55-0400

The standard error of the mean is the standard deviation of the sample means. If the population standard deviation is "\sigma" and the sample size is n, then the standard deviation of the sample means is:

"\sigma_{\bar{x}} = \frac{\sigma}{\sqrt{n}} \\

\sigma_{\bar{x}} = \frac{3.15}{\sqrt{5}} = 1.40"